# Graph-Based Bidirectional Transformer Decision Threshold Adjustment Algorithm for Class-Imbalanced Molecular Data

### v1

@@ 2.1.1 MBO Scheme with Decision Threshold Adjustment
After repeating the above steps for a specified number of iterations, our model generates a predicted label matrix for each threshold. To select the optimal decision threshold from the set of thresholds for a given labeled/unlabeled partition of a data set, our model then computes the ROC-AUC score for each label matrix and subsequently chooses the threshold whose predicted labels yield the highest ROC-AUC score. Notably, the new thresholding procedure only involves alterations to the displacement step and does not require additional iterations of the diffusion or projection to simplex steps. This allows one to save computational time, in contrast to many cost-sensitive methods which require the complete retraining of the classifier. Section 3.3 contains further discussion regarding this threshold optimization and its results.
Algorithm 1 DT-MBO Algorithm (using techniques from [43, 45])
Labeled data =$\mathscr{L}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$, where $y_{i}$ is the label of the compound whose fingerprint is $\mathbf{x}_{i}$, $N$ (# of data set elements), $N_{n}$ (# of nearest neighbors), $N_{e}$ (# of eigenvectors to be computed), $C$, $dt>0$, $N_{t}$ (maximum # of iterations), $N_{p}$ (# of labeled points or # of points in training set), $N_{l}$ (# of labeled sets), $N_{s}$ (# of times to apply diffusion operator in MBO method), $m$ (# of classes), $\tau_{\text{low}}$ (lowest threshold to test), $\tau_{\text{high}}$ (highest threshold to test), $k_{\text{min}}$ (index of the minority class).
@@
1: Construct graph with $N_{n}$ nearest neighbors or go straight to Step 3 using the Nyström technique [17, 18, 3] (useful for very large data sets).
2: Compute the graph Laplacian $\mathbf{L}$.
3: Compute (or approximate using the Nyström method) the first $N_{e}$ eigenvalues $\Lambda$ and eigenvectors $\Phi$ of $\mathbf{L}$.
4: Complete the following steps:
@@
     Randomly generate $N\times 1$ vector $\Gamma$, where $\Gamma_{j}=1$ for $N_{p}$ labeled points and $\Gamma_{j}=0$ otherwise.
@@
         $\mathbf{U}_{ik}^{~{}0}\leftarrow rand((0,1))$, except for labeled nodes, where $\mathbf{u}_{i}^{0}\leftarrow\boldsymbol{e}_{k}$ for labeled nodes, where $k$ is the true class.
         $\mathbf{u}_{i}^{0}\leftarrow projectToSimplex(\mathbf{u}_{i}^{0})$, where $\mathbf{u}_{i}$ is $i^{th}$ row of $\mathbf{U}^{0}$.
@@
              $\mathbf{B}\leftarrow\mathbf{C}(\Phi^{\prime}(\Gamma\cdot(\mathbf{U}^{n}-\mathbf{U}^{0})))$, with $\cdot$ indicating row-wise multiplication.
@@
     Compute ROC-AUC score for each threshold using its predicted labels, and choose the threshold that yields the highest score.
@@
2.1.2 Bidirectional Transformer
As in our previous work [28], the BTDT-MBO algorithm utilizes a bidirectional transformer to convert the SMILES input for each molecular compound to a latent space vector, which is then extracted as a molecular fingerprint (called a BT-FP) for that compound. After being normalized, the BT-FPs for the compounds in a given data set comprise the features passed to the DT-MBO method for that data set. Specifically, the BTDT-MBO algorithm uses a particular bidirectional transformer model introduced by Chen et al. [8], which is based on an attention mechanism for self-supervised learning (SSL). The model, the self-supervised learning platform (SSLP), enables the generation of BT-FPs without the need for data labels. Moreover, the model used for BTDT-MBO can be chosen out of three SSLP models constructed by Chen et al. [8]: one model pretrained on the ChEMBL data set [24], one model pretrained on the union of the ChEMBL and PubMed [35] data sets, and one model trained on the union of the ChEMBL, PubMed, and ZINC [30] data sets. Additonally, there is an option to fine-tune the selected SSLP for specific predictive tasks.
In our prior work [28], we demonstrated the predictive success of the BT-MBO model utilizing the self-supervised learning platform trained solely on ChEMBL for scarcely labeled molecular data. Consequently, the proposed BTDT-MBO model in the present work uses the same SSLP trained on ChEMBL. Similarly, our BT-MBO model in [28] performed well without fine-tuning the SSLP, so we also bypass the fine-tuning step for the proposed BTDT-MBO model, instead passing our data directly to the pretrained SSLP.
Algorithm 2 BTDT-MBO Algorithm (using similar techniques as in [28])
Unlabeled data $\mathscr{U}=\{s_{i}\}_{i=1}^{N}$, where $s_{i}$ is the SMILES string of a molecular compound, a selected pretrained model (from model pretrained on ChEMBL, model pretrained on union of ChEMBL and PubMed, or model pretrained on the union of ChEMBL, PubMed, and ZINC), a dictionary that assigns an integer value to each SMILES symbol.
Set of normalized fingerprints $\mathscr{F}_{\text{BT}}=\{\mathbf{x}_{i}\}_{i=1}^{N}$.
1: Preprocessing: binarize input SMILES data.
2: Load selected pretrained model.
3: Extract hidden information from $\mathscr{U}$ using the pretrained model:
Initialize empty dictionary $D_{\text{hidden}}$ of length $N$ of hidden features.
@@ 2.1.2 Bidirectional Transformer
     Generate a tensor $\mathbf{t}_{i}$ with dimension equal to the length of the string $s_{i}$, where each element of $\mathbf{t}_{i}$ is the integer dictionary value of the character in the corresponding position in $s_{i}$.
     Pass $\mathbf{t}_{i}$ to the bidirectional encoder transformer from the pretrained model and extract the inner state from the last hidden layer $\mathbf{T}_{i}$, whose dimensions are $(l_{i},1,512)$ (where $l_{i}$ is the length of $\mathbf{t}_{i}$).
     Reshape $\mathbf{T}_{i}$ into a tensor $\mathbf{V}_{i}$ with dimensions $(l_{i},512)$.
     Save $\mathbf{V}_{i}$ as $i$th element of $D_{\text{hidden}}$.
@@
4: Generate fingerprints from hidden information:
Initialize empty list of fingerprints $\mathscr{F}_{\text{BT}}$ of length $N$.
@@
     Extract first row $\mathbf{x}_{i}$ from the $i$th element of $D_{\text{hidden}}$.
     Save $\mathbf{x}_{i}$ as $i$th fingerprint in $\mathscr{F}_{\text{BT}}$.
@@
5: Scale fingerprints $\mathscr{F}_{\text{BT}}=\{\mathbf{x}_{i}\}_{i=1}^{N}$ to have zero mean and unit variance.
6: Send scaled fingerprints $\mathscr{F}_{\text{BT}}=\{\mathbf{x}_{i}\}_{i=1}^{N}$ to the DT-MBO method in Algorithm 1.
@@ 2.1.3 Overall Procedure
  • Molecular data in the form of SMILES strings are passed to the self-supervised learning platform pretrained on ChEMBL data set [24].
  • The resulting BT-FPs are scaled to have zero mean and unit variance and then passed to the DT-MBO method, outlined in Algorithm 1.
More details of the stages are given in Algorithm 2. Notably, because the SSLP does not require data labels, this procedure can be applied to data with any number of classes. Paired with our DT-MBO method as outlined in Section 2.1.1 and Algorithm 1, our BTDT-MBO model can thus be applied to problems with an arbitrary number of classes.
@@ 2.2 Distance Correlation
As discussed in Section 2.1.1, the similarity graph-based framework on which the DT-MBO scheme operates requires a weight function to measure the degree of similarity between two data elements. Various weight functions can be used, but a suitable weight function will be defined such that a high degree of similarity between two data elements is reflected by a large weight function output, and a low degree of similarity is reflected by a small weight function output [59].
As in the authors’ previous work [28] using the BT-MBO method, our proposed BTDT-MBO method utilizes a Gaussian kernel (1) as one of the ways to compute weights for the MBO scheme. By construction, when the distance between two data elements is 0, the Gaussian weight function evaluated at those two data elements equals 1. As the distance between two data elements approaches infinity, the Gaussian weight function approaches 0. In other words, the Gaussian weight function outputs values close to 0 for dissimilar data elements and 1 for identical data elements, with a weight function output closer to 1 reflecting greater similarity.
In the present work, we further propose distance correlation [57] as a potential weight function for the DT-MBO scheme, motivated in part by the success of Hozumi et al. [29] in applying distance correlation to feature clustering. Also, as will be illustrated below, the distance correlation weight function scales similarly to the Gaussian weight function, with potential outputs between 0 and 1, and higher weight function values corresponding to greater similarity, supporting our substitution. We hope to provide a new interpretation of the MBO method by incorporating the distance correlation as a weight function.
Distance correlation [57] is a measure of dependence between random vectors that generalizes of the idea of correlation. In particular, distance correlation is defined for pairs of vectors in arbitrary dimension, and the distance correlation of two vectors is 0 if and only if the two vectors are independent. We use the distance correlation between two vectors as defined in [57] and use the same notation below as in [29]. First, given a vector $\mathbf{z}^{i}$, $i=1,2,\dots I$, we can compute a distance matrix with entries defined by
$a^{i}_{jk}=||z^{i}_{m}-z^{i}_{k}||,\quad m,k=1,2,\dots M,$ (4)
where |||||$\cdot$||| | ⋅ | | is the Euclidean norm. Also, the doubly centered distance for vector $\mathbf{z}^{i}$ is given by
$A^{i}_{jk}:=a_{jk}-\bar{a}_{j\cdot}-\bar{a}_{\cdot k}+\bar{a}_{\cdot\cdot}.$ (5)
Here, $\bar{a}_{j\cdot}$ is the $j$th row mean, $\bar{a}_{\cdot k}$ is the $k$th column mean, and $\bar{a}_{\cdot\cdot}$ is the grand mean of the distance matrix for $\mathbf{z}^{i}$.
Given a pair of vectors $\mathbf{z}^{i}$ and $\mathbf{z}^{j}$, we can define the squared distance covariance as
@@
Now, the distance correlation between vectors $\mathbf{z}^{i}$ and $\mathbf{z}^{j}$ is defined as
@@
Note that the distance correlation $\text{dCor}(\mathbf{z}^{i},\mathbf{z}^{j})$ can take on values in the range $[0,1]$ (see a complete proof of this fact in [57]). We have $\text{dCor}(\mathbf{z}^{i},\mathbf{z}^{j})=0$ when vectors $\mathbf{z}^{i}$ and $\mathbf{z}^{j}$ are independent (or $\text{dCov}^{2}(\mathbf{z}^{i},\mathbf{z}^{j})=0$), and we have $\text{dCor}(\mathbf{z}^{i},\mathbf{z}^{j})=1$ when $\mathbf{z}^{i}$ and $\mathbf{z}^{j}$ are linearly dependent. Additionally, a higher squared distance covariance between two vectors corresponds to a distance correlation closer to 1 for those vectors.
Thus, the distance correlation (7) satisfies the conditions for a similarity weight function for the MBO scheme. As discussed above, the range of the distance correlation is identical to that of the Gaussian kernel, so computationally implementing the distance correlation as a weight function in our DT-MBO method is straightforward. Specifically, only the construction of the graph Laplacian $\mathbf{L}$ in Step 2 of Algorithm 1 differs when using the distance correlation to compute the graph weights. Furthermore, Székely et al. [57] proved relationships between the distance correlation and the standard bivariate normal distribution which support this substitution in our algorithm. Namely, if $X$ and $Y$ have standard normal distributions with correlation $\rho$, then $\text{dCor}(X,Y)\leq|\rho|$.
@@ 3.1 Data Sets
The data sets used in the present work are characterized by a high level of imbalance in class sizes, measured by the imbalance ratio (IR). The IR of a data set with two classes is defined as the ratio of the number of points in the majority class to the number of points in the minority class [12]. Out of the public data sets Esposito et al. [12] used for benchmarking their proposed thresholding models, we selected data sets from each of the following classes to benchmark our methods:
  • DS1 Data Sets (obtained from [12], introduced in [50] and refined in [51]): Each of the DS1 data sets corresponds to a particular ChEMBL target and contains 100 associated diverse active compounds against that target. For each data set, we randomly chose 2,000 inactives from a set of 10,000 assumed inactives drawn from ZINC [31, 55], reproducing the data set construction procedure from [12]. These 10,000 compounds were selected in [51] to have similar property distributions to the set of actives, with similarity measured by an atom-count fingerprint. Every DS1 data set has an IR of 20.0 by construction.
  • DrugMatrix Assays (obtained from [12]): Each the 44 DrugMatrix assays benchmarked in [12] contains results for 842 compounds tested in a particular assay, with each data set labeled by the assay tested. Each molecule was labeled as active or inactive based on its ”activity_comment” value in the source data. In the present work, we test our proposed method on four of these data sets, with IRs ranging from 16.5 to 20.0. The source data was retrieved from ChEMBL and was originally recorded in the DrugMatrix database [1].
A brief summary of the data sets used for testing the proposed BTDT-MBO model is given in Table 1.
The authors’ BT-MBO method from [28] performed better than state-of-the-art methods for scarcely labeled molecular data. Inspired by this success, in the present work, we focus on the challenging task of making predictions on highly imbalanced data sets. All data sets used for benchmarking the proposed BTDT-MBO method have an IR of 16.5 or greater.
@@
CHEMBL1909150 [12, 1] DrugMatrix 842 16.9 Active, Inactive
CHEMBL1909157 [12, 1] DrugMatrix 842 18.6 Active, Inactive
CHEMBL1909132 [12, 1] DrugMatrix 842 20.0 Active, Inactive
CHEMBL1909134 [12, 1] DrugMatrix 842 16.5 Active, Inactive
CHEMBL100 [12, 50, 51] DS1 2100 20.0 Active, Inactive
CHEMBL8 [12, 50, 51] DS1 2100 20.0 Active, Inactive
Table 1: Information about the data sets used for benchmarking the model (results for these data sets are shown in Figures 2 and 3). All data sets group molecules into two classes.

### v3

@@ 2.1.1 MBO Scheme with Decision Threshold Adjustment
After repeating the above steps for a number of iterations, our model generates a predicted label matrix for each threshold. To select the optimal decision threshold from the set of thresholds for a given labeled/unlabeled partition of a data set, our model then computes the ROC-AUC score for each label matrix and subsequently chooses the threshold whose predicted labels yield the highest ROC-AUC score. Notably, the new thresholding procedure only involves alterations to the displacement step and does not require additional iterations of the diffusion or projection to simplex steps, thus saving computational time, in contrast to many cost-sensitive methods which require the complete retraining of the classifier. Section 3.3 contains discussion regarding this threshold optimization and its results.
2.1.2 Bidirectional Transformer
As in our previous work 40, the BTDT-MBO algorithm utilizes a bidirectional transformer to convert the SMILES input for each molecular compound to a vector in a latent space, which is then obtained as a molecular fingerprint, named BT-FP, for that compound. After being normalized, the BT-FPs for the compounds in a given data set comprise the features passed to the DT-MBO method for that data set. Specifically, the BTDT-MBO algorithm uses a particular bidirectional transformer model introduced by Chen et al. 43 involving an attention structure for self-supervised learning (SSL). The model, the self-supervised learning platform (SSLP), enables the generation of BT-FPs without the need for data labels. Moreover, the model used for BTDT-MBO can be chosen out of three SSLP models constructed by Chen et al. 43. The three SSLP models include one pretrained on the ChEMBL data set 58. The other two models include one pretrained on the union of the ChEMBL and PubMed 59 data sets and one model trained on the union of the ChEMBL, PubMed, and ZINC 60 data sets. Additionally, there is an option to fine-tune the selected self-supervised learning platform for specific predictive tasks.
In our prior work 40, we demonstrated the predictive success of the BT-MBO model utilizing the self-supervised learning platform trained solely on ChEMBL for scarcely labeled molecular data. Consequently, the proposed BTDT-MBO model in the present work uses the same SSLP trained on ChEMBL. Similarly, our BT-MBO model in 40 performed well without fine-tuning the SSLP, so we also bypass the fine-tuning step for the proposed BTDT-MBO model, instead passing our data directly to the pretrained SSLP.
Algorithm 1 DT-MBO Algorithm (using techniques from 56, 49)
Labeled set =$\mathscr{L}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$, with fingerprint $\mathbf{x}_{i}$ and label $y_{i}$, $N$= size of data set, $N_{n}=$number of nearest neighbors, $N_{e}=$number of eigenvectors, $dt$, $C$, $N_{t}=$ maximum number of iterations, $N_{p}=$ number of labeled elements, $N_{l}=$ number of labeled collections, $N_{s}$, $m=$ number of classes, $\tau_{\text{low}}=$ lowest threshold, $\tau_{\text{high}}=$ highest threshold, $k_{\text{min}}=$ index of minority group/cluster.
@@ 2.1.2 Bidirectional Transformer
1: Compute the $N_{n}$-nearest neighbor graph, unless one is using the Nyström technique 61, 62, 63 for large data. In the latter case, go to Step 3.
2: Construct the graph Laplacian $\mathbf{L}$.
3: Calculate the smallest $N_{e}$ eigenvalues $\Lambda$ and the corresponding eigenvectors $\Phi$ of $\mathbf{L}$. One can also approximate them using the Nyström method 61, 62, 63 in case of large data sets.
4: The main procedure of the proposed method:
@@
     Let the $N\times 1$ vector $\Gamma$ vector be defined so that $\Gamma_{j}=1$ for labeled points and 00 for unlabeled points.
@@
         Initialize each value of a row of $\mathbf{U}$ to a random number in [0,1], except for rows corresponding to labeled nodes, in which case, set $\mathbf{u}_{i}^{0}\leftarrow\boldsymbol{e}_{k}$, where $k$ is the true class.
         $\mathbf{u}_{i}^{0}\leftarrow projectToSimplex(\mathbf{u}_{i}^{0})$. Here, $\mathbf{u}_{i}$ is $i^{th}$ row of the matrix $\mathbf{U}^{0}$.
@@
              $\mathbf{B}\leftarrow\mathbf{C}(\Phi^{\prime}(\Gamma\cdot(\mathbf{U}^{n}-\mathbf{U}^{0})))$, where row-wise multiplication is performed.
@@
     Compute ROC-AUC score for each threshold, and choose the threshold that yields the highest score.
@@
Algorithm 2 BTDT-MBO Method (using related procedures as in 40)
Unlabeled set $\mathscr{U}=\{s_{i}\}_{i=1}^{N}$, where $s_{i}$ is a molecular compound’s SMILES string), a chosen pre-trained model from the three described in Section 2.1.2, a dictionary which designates an integer number to each SMILES element.
$\mathscr{F}_{\text{BT}}=\{\mathbf{x}_{i}\}_{i=1}^{N}$, a set of normalized fingerprints.
1: Binarize the data consisting of SMILES strings.
2: Loading step: load the chosen model (which is pre-trained).
3: Extracting step: Extract the hidden information from $\mathscr{U}$ via the chosen pretrained model:
Initialize the empty dictionary $D_{\text{hidden}}$ (of length $N$) containing hidden features.
@@
     Compute tensor $\mathbf{t}_{i}$ (of the same dimension as $s_{i}$), containing the dictionary values of the characters in the analogous locations in $s_{i}$.
     Let $\mathbf{t}_{i}$ be an input to the bidirectional encoder transformer from the chosen pretrained model. Obtain the inner state from the last hidden layer $\mathbf{T}_{i}$, whose dimensions are $(l_{i},1,512)$ (where $l_{i}$ is the length of $\mathbf{t}_{i}$).
     Reshape $\mathbf{T}_{i}$ into a tensor $\mathbf{V}_{i}$ with the following dimensions: $(l_{i},512)$.
     Set $\mathbf{V}_{i}$ as $i$th element of the matrix $D_{\text{hidden}}$.
@@
4: Generating step: Produce the fingerprints from the hidden information:
Initialize the list of fingerprints $\mathscr{F}_{\text{BT}}$ (of length $N$).
@@
     Compute the first row $\mathbf{x}_{i}$ from the $i$th element of the matrix $D_{\text{hidden}}$.
[…]
@@
5: Scale fingerprints $\mathscr{F}_{\text{BT}}=\{\mathbf{x}_{i}\}_{i=1}^{N}$ so that they have unit variance and zero mean.
6: Send scaled fingerprints $\mathscr{F}_{\text{BT}}=\{\mathbf{x}_{i}\}_{i=1}^{N}$ to the DT-MBO algorithm, described in Algorithm 1.
@@ 2.1.3 Overall Procedure
  • Molecular data in the form of SMILES strings are passed to the self-supervised learning platform pretrained on ChEMBL data set 58.
  • The resulting BT-FPs are the rescaled so that they have zero mean as well as unit variance and then served as an input to the DT-MBO method, outlined in Algorithm 1.
More details of the stages are given in Algorithm 2. Notably, because the SSLP does not require data labels, this procedure can be applied to data with any number of classes. Paired with our DT-MBO method as outlined in Section 2.1.1 and Algorithm 1, our BTDT-MBO model can thus be used for problems with any number of clusters or classes.
@@ 2.2 Distance Correlation
As discussed in Section 2.1.1, the similarity graph-based framework on which the DT-MBO scheme operates requires a weight function to compute the degree of similarity between two data elements. Various weight functions can be used, but a suitable weight function is defined so that a high degree of similarity between two elements is reflected by a large weight function output, and a low degree of similarity is reflected by a small weight function output 64.
As in the authors’ previous work 40 using the BT-MBO method, our proposed BTDT-MBO method utilizes a Gaussian kernel (1) as one of the ways to compute weights for the MBO scheme. By construction, when the distance between two data elements is 0, the Gaussian weight function evaluated at those two data elements equals 1. As the distance between two data elements approaches infinity, the Gaussian weight function approaches 0. In other words, the Gaussian weight function outputs values close to 0 for dissimilar data elements and 1 for identical data elements, with a weight function output closer to 1 reflecting greater similarity.
In the present work, we further propose distance correlation 46 as a potential weight function for the DT-MBO scheme, motivated in part by the success of Hozumi et al. 47 in applying distance correlation to feature clustering. Also, as will be illustrated below, the distance correlation weight function scales similarly to the Gaussian weight function, with potential outputs between 0 and 1, and higher weight function values corresponding to greater similarity, supporting our substitution. We hope to provide a new interpretation of the MBO method by incorporating the distance correlation as a weight function.
Distance correlation 46 is a measure of dependence between random vectors that generalizes of the idea of correlation. In particular, distance correlation is defined for pairs of vectors in arbitrary dimension, and the distance correlation of two vectors is 0 if and only if the two vectors are independent. We use the distance correlation between two vectors as defined in 46 and use the same notation below as in 47. First, given a vector $\mathbf{z}^{i}$, $i=1,2,\dots I$, we can compute a distance matrix with entries defined by
$a^{i}_{jk}=||z^{i}_{m}-z^{i}_{k}||,\quad m,k=1,2,\dots M.$ (4)
Here, |||||$\cdot$||| | ⋅ | | represents the Euclidean norm. Also, the doubly centered distance for $\mathbf{z}^{i}$ is given by
$A^{i}_{jk}:=a_{jk}-\bar{a}_{j\cdot}-\bar{a}_{\cdot k}+\bar{a}_{\cdot\cdot},$ (5)
where $\bar{a}_{j\cdot}$ represents the $j$th row mean. In addition, $\bar{a}_{\cdot k}$ represents the $k$th column mean, and $\bar{a}_{\cdot\cdot}$ represents the grand mean of the distance matrix for $\mathbf{z}^{i}$.
Given $\mathbf{z}^{i}$ and $\mathbf{z}^{j}$, we can define the squared distance covariance as
@@
Now, the distance correlation between $\mathbf{z}^{i}$ and $\mathbf{z}^{j}$ is defined as
@@
Note that the distance correlation $\text{dCor}(\mathbf{z}^{i},\mathbf{z}^{j})$ can take on values in the range $[0,1]$ (see a complete proof of this fact in 46). We have $\text{dCor}(\mathbf{z}^{i},\mathbf{z}^{j})=0$ when vectors $\mathbf{z}^{i}$ and $\mathbf{z}^{j}$ are independent (or $\text{dCov}^{2}(\mathbf{z}^{i},\mathbf{z}^{j})=0$), and we have $\text{dCor}(\mathbf{z}^{i},\mathbf{z}^{j})=1$ when $\mathbf{z}^{i}$ and $\mathbf{z}^{j}$ are linearly dependent. Additionally, a higher squared distance covariance between two vectors corresponds to a distance correlation closer to 1 for those vectors.
Thus, the distance correlation (7) satisfies the conditions for a similarity weight function for the MBO scheme. As discussed above, the range of the distance correlation is identical to that of the Gaussian kernel, so computationally implementing the distance correlation as a weight function in our DT-MBO method is straightforward. Specifically, only the construction of the graph Laplacian $\mathbf{L}$ in Step 2 of Algorithm 1 differs when using the distance correlation to compute the graph weights. Furthermore, Székely et al. 46 proved relationships between the distance correlation and the standard bivariate normal distribution which support this substitution in our algorithm. Namely, if $X$ and $Y$ have standard normal distributions with correlation $\rho$, then $\text{dCor}(X,Y)\leq|\rho|$.
@@ 3 Results and Discussion
In this section, we discuss the computational experiments and results. The code used for the DT-MBO algorithm was written in MATLAB. To carry out computational experiments and optimize parameters, we utilized the High Performance Computing Center (HPCC) at Michigan State University. For all jobs submitted through HPCC, we specified two CPUs per task. Distance correlation weights were computed using the distance_correlation function from the dcor library in Python.
@@ 3.1 Data Sets
The data sets used in the present work are characterized by a high level of imbalance in class sizes, measured by the imbalance ratio (IR). The IR of a data set with two classes is defined as the ratio of the number of points in the majority class to the number of points in the minority class 1. Out of the public data sets Esposito et al. 1 used for benchmarking their proposed thresholding models, we selected data sets from each of the following classes to benchmark our methods:
  • DS1 Data Sets (obtained from 1, introduced in 65 and refined in 66): Each of the DS1 data sets corresponds to a particular ChEMBL target and contains 100 associated diverse active compounds against that target. For each data set, we randomly chose 2,000 inactives from a set of 10,000 assumed inactives drawn from ZINC 67, 68, reproducing the data set construction procedure from 1. These 10,000 compounds were selected in 66 to have similar property distributions to the set of actives, with similarity measured by an atom-count fingerprint. Every DS1 data set has an IR of 20.0 by construction.
  • DrugMatrix Assays (obtained from 1): Each the 44 DrugMatrix assays benchmarked in 1 contains results for 842 compounds tested in a particular assay, with each data set labeled by the assay tested. Each molecule was labeled as active or inactive based on its ”activity_comment” value in the source data. In this paper, we evaluate our proposed procedure on four of these data sets, with IRs ranging from 16.5 to 20.0. The source data was retrieved from ChEMBL and was originally recorded in the DrugMatrix database 69.
A brief summary of the data sets used for testing the BTDT-MBO model is given in Table 1.
The authors’ BT-MBO method from 40 performed better than state-of-the-art techniques for scarcely labeled molecular data. Inspired by this success, in the present work, we focus on the challenging task of making predictions on highly imbalanced data sets. All data sets used for benchmarking the proposed BTDT-MBO method have an IR of 16.5 or greater.
@@
CHEMBL1909150 DrugMatrix 842 16.9 Active, Inactive
CHEMBL1909157 DrugMatrix 842 18.6 Active, Inactive
CHEMBL1909132 DrugMatrix 842 20.0 Active, Inactive
CHEMBL1909134 DrugMatrix 842 16.5 Active, Inactive
CHEMBL100 DS1 2100 20.0 Active, Inactive
CHEMBL8 DS1 2100 20.0 Active, Inactive
Table 1: Information about the data used for benchmarking the model (results are shown in Figures 3 and 7). All data sets group molecules into two classes. The first four data sets are from 1, 69 and the last two data sets are from 1, 65, 66.